OpenStudy (anonymous):

Let g(x)=x^2+2x-6. Find the value of 2[g(2)]?

OpenStudy (whpalmer4):

If \(g(x) = x^2+2x-6\), what is the value of \(g(2)\)? Replace all the \(x\)'s by 2 and compute the result.

OpenStudy (whpalmer4):

Multiply that by 2 to get the value of 2[g(2)]... What do you get?

OpenStudy (anonymous):

4?

OpenStudy (whpalmer4):

No, I think you meant to type "4!" :-) Yes, 4 is correct.

OpenStudy (anonymous):

Haha thanks :)

OpenStudy (whpalmer4):

Not so hard, right?

OpenStudy (anonymous):

Just the 2[] threw me off :p

OpenStudy (whpalmer4):

Here's a slightly trickier one: let \(f(x) = x+3\) and \(g(x) = x^2-5\) what is the value of \(g(f(x))\) if \(x = 3\)? Sometime you see this written as \((g\circ f)(x)\)